# HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation

### v2

@@ 4.2 Main Results
We evaluate our approach on two hypergraph tasks: node classification and hyperedge prediction (see Appendix B.1). Node classification evaluation. We first evaluate HiTeC on the node classification task. Table 2 presents the results, revealing two key observations: (1) Graph-based SSL methods perform consistently worse than HSSL methods. This is primarily due to the loss of high-order relations caused by clique expansion. Even GIANT, which incorporates textual and structural signals, remains insufficient for modeling high-order structures. These results highlight the necessity of developing SSL methods tailored for TAHG. (2) HiTeC achieves the best overall performance with strong scalability. VilLain performs poorly due to its neglect of textual information, highlighting the critical role of node semantics. TriCL and SE-HSSL also underperform, mainly because of inadequate textual feature extraction and limited modeling of long-range dependencies. While both HyperBERT and HiTeC integrate textual semantics with high-order structure, HyperBERT fails to fully exploit semantic–structural signals and cannot scale to even medium-sized hypergraphs (e.g., History, Photo) due to its joint modeling scheme. In contrast, HiTeC consistently outperforms all baselines, surpassing the second-best method by 9.77% and 10.26% on Cora and Fitness, respectively. This is attributed to the semantic-aware augmentations and $s$-walk-based subgraph contrast, which reduce perturbation noise and enable the model to capture more global dependencies. Furthermore, our two-stage decoupled design brings scalability, with the structure-aware contrastive objective complementing it while preserving high performance.
(a) Impact of components.
(b) Impact of prompt.
Figure 4: Ablation study. (a) Impact of components: “w/o pre” removes the structure-aware pre-training of the text encoder; “w/o shd” replaces the semantic-aware hyperedge drop with the random drop; and “w/o s” removes the subgraph-level contrastive objective. (b) Impact of prompt: “w/o prompt” uses original text only; “w/o domain”, “w/o topology”, and “w/o context” remove the corresponding content from the prompt template.
@@ 4.3 Ablation Study
We conduct ablation studies on: (1) key components of HiTeC, (2) prompt template elements, and (3) text encoder backbones (see Appendix B.2).
Impact of components. We first assess the contribution of each component in HiTeC. As shown in Figure 4(a), removing any module results in a performance drop. The most severe drop is caused by “w/o pre”, with 11.25% and 23.41% declines on Computers and Fitness, respectively. This highlights the importance of incorporating topology into the text encoder for learning expressive textual representations. Semantic-aware hyperedge drop is also critical, as its removal (“w/o shd”) leads to a 17.02% drop on Fitness, demonstrating the value of maintaining semantic consistency during structural perturbation.
Impact of prompt. We further analyze the effect of prompt design in Figure 4(b). Overall, our prompt-enhanced augmentation outperforms other variants. “w/o prompt” yields the largest performance drop, confirming the benefits of injecting diverse semantics into text augmentation. Removing individual content, such as context or topology, also leads to noticeable declines, highlighting the importance of incorporating meaningful structural signals into the original text.
@@ 4.4 Parameter Sensitivity Analysis
We conduct sensitivity experiments on two key hyperparameters: the overlap size $s$ in the $s$-walk, and the subgraph sampling ratio $r$ (see Appendix B.3).
Figure 5: Sensitive analysis of $s$. The line plot shows model accuracy under different values of $s$. The bar plot indicates the average number of hyperedges in the sampled subgraph, reflecting its exploration range.
Figure 6: Training time vs. accuracy of random walk and $s$-walk. Bars indicate training time, markers denote accuracy.
Overlap size $s$. Figure 5 illustrates how model performance varies with different values of $s$, where model performance first improves and then degrades as $s$ increases. This trend can be attributed to the trade-off between structural coherence and exploration: when the exploration range remains sufficiently broad, a larger $s$ promotes more cohesive subgraphs while preserving essential contextual semantics. However, excessively strict overlap constraints (e.g., ss=10) limit structural diversity and impair performance. Figure 6 further indicates that increasing $s$ leads to smaller subgraphs, thereby accelerating training. These results highlight the importance of selecting an appropriate $s$ to balance accuracy and efficiency.
@@ 5 Related Work
Our work is closely related to self-supervised learning on hypergraphs and representation learning on text-attributed graphs (see Appendix C.1).
@@ 5.1 Self-supervised Learning on Hypergraphs
SSL provides a promising alternative for hypergraph representation learning without expensive labels. Most existing HSSL methods focus on contrastive learning. HyperGCL (Wei et al. 2022) introduces a learnable augmentation strategy, while TriCL (Lee and Shin 2023) and SE-HSSL (Li et al. 2024) design hierarchical objectives across multiple granularities. Beyond contrastive approaches, generative SSL has also emerged, with HypeBoy (Kim et al. 2024a) proposing a hyperedge-filling task. Other efforts address different settings, such as VilLain (Lee, Lee, and Shin 2024), which targets feature-absent scenarios. While effective in modeling high-order structure, these methods overlook the rich textual semantics in TAHGs. HyperBERT (Bazaga, Lio, and Micklem 2024) pioneers SSL on TAHGs by jointly encoding textual semantics and hypergraph structure, but it incurs high computational overhead and suffers from poor scalability. To bridge the gap, we explore a scalable and effective contrastive learning framework for TAHGs.
@@ 6 Conclusion
We propose HiTeC, a hierarchical contrastive learning framework with semantic-aware enhancement, to address limitations of existing HCL methods in TAHGs. To ensure scalability, HiTeC adopts a two-stage decoupled paradigm. The first stage pre-trains a text encoder with a structure-aware contrastive objective, alleviating the graph-agnostic nature of conventional encoders. In the second stage, two semantic-aware augmentation strategies, including prompt-enhanced text augmentation and semantic-aware hyperedge drop, are designed to generate more informative hypergraph views. Furthermore, we introduce a hierarchical contrastive objective, where a novel $s$-walk-based subgraph-level contrast captures long-range structural dependencies beyond node and hyperedge scopes. Extensive experiments on six real-world datasets and multiple downstream tasks demonstrate the effectiveness and scalability of HiTeC.
@@
Appendix A Additional Experimental Setup
A.1 Datasets
@@
Co-citation hypergraphs. We construct two academic hypergraphs, Citeseer (Giles, Bollacker, and Lawrence 1998) and Cora (Sen et al. 2008), based on the standard citation networks, where each node represents a scientific publication. The textual attributes of each node, including the title and abstract, are directly adopted from prior works (He et al. 2024; Wang et al. 2024). To capture high-order relationships, we define a hyperedge as the set of all papers that co-cite the same reference.
@@
E-commerce hypergraphs. We further construct four e-commerce hypergraphs from the Amazon product network (Ni, Li, and McAuley 2019): Book-History, Ele-Photo, Ele-Computers, and Sports-Fitness. In these datasets, each node represents a product, and edges reflect frequent co-purchase or co-browsing behavior. The raw textual attributes of nodes are obtained from the annotations provided by (Yan et al. 2023). Specifically, in Book-History, textual features include book titles and descriptions; in Ele-Computers and Ele-Photo, they consist of high-rated reviews and product summaries; in Sports-Fitness, only product titles are used. To induce high-order structures, we adopt a max-clique-based reconstruction strategy (Bron and Kerbosch 1973; Wang and Kleinberg 2024), where hyperedges are formed by identifying maximal cliques in the co-purchase graph.
A.2 Baselines
@@
We compare our method against seven strong self-supervised learning (SSL) baselines, covering both graph-based and hypergraph-based paradigms.
@@
Graph-based SSL. In this work, we adopt three widely used graph-based SSL frameworks: two contrastive learning methods, GraphCL (You et al. 2020) and BGRL (Thakoor et al. 2022), and one generative method, GraphMAE (Hou et al. 2022). To adapt these methods to hypergraphs, we apply clique expansion following the protocol in (Lee and Shin 2023). For textual feature extraction, we employ shallow encoders (SE) implemented via Skip-Gram (Mikolov et al. 2013), as well as GIANT (Chien et al. 2022a), a graph-agnostic text encoder designed for raw textual attributes.
@@
Hypergraph-based SSL. We include four state-of-the-art hypergraph SSL methods: two contrastive approaches, TriCL (Lee and Shin 2023) and SE-HSSL (Li et al. 2024); one generative method, HypeBoy (Kim et al. 2024a); and VilLain (Lee, Lee, and Shin 2024), which targets the feature-absent setting. All methods adopt the same SE setup and are additionally evaluated with two pre-trained language models: BERT-Base (Devlin et al. 2019) and RoBERTa-Base (Liu et al. 2019).
A.3 Evaluation Protocol
@@
We evaluate our method on two standard hypergraph learning tasks: node classification and hyperedge prediction. After pretraining with HiTeC, we adopt a linear evaluation protocol (Lee and Shin 2023; Li et al. 2024; Kim et al. 2024a), where the pretrained node embeddings are frozen and used as input features for downstream classifiers.
@@
Node classification. Following (Lee and Shin 2023; Li et al. 2024), we randomly split each dataset into 10%/10%/80% for training, validation, and testing. Then, a simple linear classifier is trained on top of the fixed node embeddings using $\ell_{2}$-regularized logistic regression. We report the average test accuracy and standard deviation over 20 random splits, each with 5 different initializations.
@@
Hyperedge prediction. Following (Kim et al. 2024a), we need to generate negative hyperedge samples to evaluate a model on the hyperedge prediction task. In our experiments, we adopt Clique Negative Sampling (CNS) (Yu et al. 2025), a widely used strategy for constructing negative hyperedge samples. Specifically, for a given hyperedge $e$, we replace a node $u\in e$ with a node $v\notin e$ such that $v$ is connected to all other nodes in $e$, i.e., $(e\setminus\{u\})\cup\{v\}$. Then, a two-layer MLP is trained on the fixed node embeddings using the 60%/20%/20% data split. Same as the node classification task, we report the average test accuracy and standard deviation over 20 random splits, each with 5 initializations.
A.4 Implementation Details
@@
All experiments are conducted on a single NVIDIA A5000 GPU (24GB memory). For all baselines, we adopt their official implementations with default settings. For HiTeC, we adopt BERT-Base (Devlin et al. 2019) as the text encoder and apply LoRA (Hu et al. 2022) for parameter-efficient fine-tuning. The text encoder is trained for 5 epochs on small datasets (Citeseer, Cora), 4 epochs on medium datasets (History, Photo), and 3 epochs on large datasets (Computers, Fitness), using a learning rate of 2e-5. The hypergraph encoder is implemented as a single-layer module. We tune the contrastive loss weight $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$. The overlap threshold $s$ and subgraph sampling ratio $r\%$ are tuned from {1, 2, 3, 4, 5, 10} and {10%, 20%, 30%, 40%, 50%}, respectively. To ensure a fair comparison, we tune the learning rate from {0.01, 0.001, 1e-4, 5e-4} and weight decay from {0, 1e-5} for all methods.
Figure 7: Training time and accuracy under different sampling ratios $r\%$. Bar plots represent training time (in minutes), while line plots indicate the corresponding accuracy. We highlight the relative speedup $\Delta t$ of $r\%{=}30\%$ over $r\%{=}50\%$. All main results are reported with $r\%{=}30\%$.

### v3

@@ 4.2 Main Results
In this section, we evaluate our approach on two representative hypergraph tasks: node classification and hyperedge prediction (see Appendix D.1).
Node classification evaluation. Table 2 presents the node classification results, revealing two key observations: (1) Graph-based SSL methods consistently underperform HSSL methods due to the loss of high-order relations caused by clique expansion. Even the TAG encoder GIANT, which jointly encodes text and pairwise structure, remains insufficient for modeling higher-order structures, highlighting the need for dedicated SSL methods for TAHGs. (2) HiTeC achieves the best overall performance across all datasets. Methods that neglect rich node features (e.g., VilLain) achieve limited performance, highlighting the importance of textual semantics in TAHG learning. HCL methods (e.g., TriCL and SE-HSSL) rely on structure-driven contrastive signals without semantic-aware augmentation, resulting in suboptimal performance. While HyperBERT jointly models textual and hypergraph information, it lacks multi-scale structural and semantic contrastive signals. In contrast, HiTeC consistently outperforms all baselines, surpassing the second-best method by 9.77% and 10.26% on Cora and Fitness, respectively. These improvements are mainly attributed to semantic-aware augmentations, which reduce perturbation noise, and the $s$-walk-based subgraph signals, which capture more informative high-order dependencies. Additional analysis on efficiency is provided in Appendix D.5.
@@ 4.3 Ablation Study
(a) Key component ablation.
(b) Text view ablation.
Figure 3: Ablation study. (a) Key component ablation: “w/o pre” removes structure-aware text encoder pre-training; “w/o shd” replaces semantic-aware hyperedge dropping with random drop; “w/o sct” removes the structure-contextualized text view; and “w/o s” removes the subgraph-level objective. (b) Text view ablation: “w/o DK”, “w/o GT”, and “w/o NC” remove domain knowledge, global topology, and neighbor context from the structure-contextualized text view, respectively.
In this section, we conduct ablation studies on: (1) key components of HiTeC, (2) structure-contextualized text view components.
Impact of key components. We evaluate the contribution of key components in HiTeC. As shown in Figure 3(a), removing any module noticeably degrades performance. The most significant drop is observed when removing text encoder pre-training (‘w/o pre”), resulting in a decrease of 13.41% on Fitness, highlighting the effectiveness of injecting topological information into the text encoder. The subgraph-level objective also plays a critical role: removing it (‘w/o s”) causes a decrease of 7.14% on Photo, demonstrating the benefit of modeling long-range dependencies inherent in TAHGs.
Impact of structure-contextualized text view components. We further evaluate the effectiveness of different components in the structure-contextualized text view in Figure 3(b). Overall, the complete view consistently outperforms all alternative variants. Specifically, removing neighbor context (“w/o NC”) and global topology (“w/o GT”) leads to pronounced performance degradation, underscoring the importance of incorporating meaningful structural signals into the original text.
@@ 4.4 Parameter Sensitivity Analysis
We investigate the impact of three key hyperparameters: the overlap size $s$, the subgraph sampling ratio $r$, and the walk length $l$ (see Appendix D.3).
[…]
Overlap size $s$. Figure 4 illustrates how model performance varies with different $s$ while fixing $l=4$. Model accuracy first improves and then degrades as $s$ increases, reflecting a trade-off between structural coherence and exploration range. Moderate overlap constraints (e.g., $s=3$) encourage cohesive subgraphs while preserving contextual diversity, leading to the best performance in most cases. In contrast, $s=1$ reduces to a vanilla random walk with weak structural correlation, while overly large overlap constraints (e.g., $s=10$) greatly restrict candidate hyperedges, reducing contextual diversity and impairing performance.
Subgraph sampling ratio $r$.
Figure 5: Training time and accuracy under different sampling ratios $r$. $\Delta t$ indicates the relative speedup.
[…]
@@ 5 Related Work
Our work is closely related to self-supervised learning on hypergraphs and representation learning on text-attributed graphs (see Appendix E.1).
@@ 5.1 Self-supervised Learning on Hypergraphs
SSL has emerged as a promising paradigm for hypergraph representation learning without expensive labels. Most existing HSSL methods are based on contrastive learning. HyperGCL Wei et al. (2022) introduces a learnable augmentation strategy, while TriCL Lee and Shin (2023) and SE-HSSL Li et al. (2024) design hierarchical contrastive objectives across multiple granularities. Beyond contrastive methods, generative SSL has also been explored, with HypeBoy Kim et al. (2024a) proposing a hyperedge-filling task. Other efforts explore alternative settings, such as feature-absent scenarios in VilLain Lee et al. (2024). While effective for modeling hypergraphs, these methods largely overlook the rich textual semantics in TAHGs, resulting in suboptimal performance. HyperBERT Bazaga et al. (2024) pioneers SSL on TAHGs by jointly optimizing text and hypergraph encoders through contrastive learning. However, its contrastive signals are limited to 1-hop hyperedge relations, restricting the modeling of long-range dependencies.
@@ 6 Conclusion
We propose HiTeC, an effective contrastive learning framework tailored for TAHGs. HiTeC adopts a two-stage paradigm that first injects topological information into text representations and then performs semantic-aware hierarchical contrastive learning on TAHGs. Specifically, two semantic-aware augmentation strategies are introduced to reduce perturbation noise, while an $s$-walk-based subgraph-level contrast captures long-range dependencies beyond node- and hyperedge-level scopes. Extensive experiments on six real-world datasets and multiple downstream tasks demonstrate the effectiveness and robustness of HiTeC.
Limitations
There are two limitations in our work: (1) The construction of the structure-contextualized text view relies on a handcrafted template. Although the proposed template demonstrates strong empirical effectiveness across multiple datasets, its optimal form may vary across different domains and tasks. Future work may explore more adaptive text-level augmentation. (2) The structure-aware pre-training stage introduces additional training overhead compared with conventional HCL methods. Nevertheless, HiTeC remains more efficient than prior TAHG methods such as HyperBERT due to its decoupled training paradigm (see Appendix D.5).
@@
Appendix A Additional Template Details
Component Description Template
[DK] Domain Knowledge You are in an Academic / E-commerce / … hypergraph. Each node represents a Paper / Product / …, and hyperedges capture Co-citation / Co-purchasing / … relationships among groups of nodes.
[GT] Global Topology Your node degree is value, indicating a Sparse / Dense connectivity pattern. You participate in the following hyperedges: Hyperedge IDs.
$t$ Raw Text Your title is Title. Your abstract or description is Abstract / Description.
[NC] Neighbor Context Below are brief descriptions of your neighboring nodes. For neighbor 1, the title is: Neighbor 1 Title and the abstract or description is :Neighbor 1 Abstract / Description. For neighbor 2 …
Table 3: Structure-contextualized text view template.
As shown in Table 3, the structure-contextualized view follows a fixed template that concatenates contextual components into a single sequence: $\tilde{t}_{1,i}=[\texttt{[DK]}\,||\,\texttt{[GT]}\,||\,\texttt{[NC]}\,||\,t_{i}].$ All components are deterministically constructed from the hypergraph structure and node attributes, without external knowledge. The number of neighboring nodes in [NC] is limited to 10 to control sequence length and ensure stable training. This design enriches structural and semantic context while preserving semantic consistency.
Appendix B Complexity Analysis
We analyze the computational complexity of the forward phase of HiTeC during training.
Text encoder pretraining. Let $l_{p}$ denote the number of layers in the PLM, $\bar{L}$ the average token length, and $d$ the hidden dimension. The PLM encoding costs $\mathcal{O}\!\left(l_{p}|\mathcal{V}|(\bar{L}d^{2}+\bar{L}^{2}d)\right)$. The triplet margin objective costs $\mathcal{O}(|\mathcal{V}|\,d)$. Therefore, the overall complexity is $\mathcal{O}\!\left(l_{p}|\mathcal{V}|(\bar{L}d^{2}+\bar{L}^{2}d)+|\mathcal{V}|d\right)$.
Hypergraph encoder pretraining. Let $l_{h}$ be the number of HNN layers and $\delta_{e}$ the average hyperedge size. The hypergraph propagation costs $\mathcal{O}\!\left(l_{h}(|\mathcal{E}|\delta_{e}d+|\mathcal{V}|d^{2})\right)$. For contrastive learning with $m$ sampled negatives (a constant), the node-, hyperedge-, and subgraph-level objectives incur costs of $\mathcal{O}(m|\mathcal{V}|d)$, $\mathcal{O}(m|\mathcal{E}|d)$, and $\mathcal{O}(mKd)$, respectively, where $K$ denotes the number of sampled subgraphs. Thus, the total complexity is $\mathcal{O}(l_{h}(|\mathcal{E}|\delta_{e}d+|\mathcal{V}|d^{2})+m(|\mathcal{V}|+|\mathcal{E}|+K)d)$.
Appendix C Additional Experimental Setup
C.1 Datasets
@@
Co-citation hypergraphs. We construct two academic hypergraphs, Citeseer Giles et al. (1998) and Cora Sen et al. (2008), from standard citation networks, where each node represents a scientific publication. Node textual attributes, including titles and abstracts, are adopted from prior works He et al. (2024); Wang et al. (2024). To capture high-order relations, we define each hyperedge as the set of papers co-citing the same reference.
@@
E-commerce hypergraphs. We further construct four e-commerce hypergraphs from the Amazon product network Ni et al. (2019): Book-History, Ele-Photo, Ele-Computers, and Sports-Fitness. Each node represents a product, while edges reflect frequent co-purchase or co-browsing behavior. Node textual attributes are adopted from Yan et al. (2023). Specifically, Book-History uses book titles and descriptions; Ele-Computers and Ele-Photo use high-rated reviews and product summaries; Sports-Fitness only uses product titles. To induce high-order structures, we adopt a max-clique-based reconstruction strategy Bron and Kerbosch (1973); Wang and Kleinberg (2024), where hyperedges are formed from maximal cliques in the original graph.
C.2 Baselines
@@
We compare our method against eight representative self-supervised learning (SSL) baselines, covering graph-based, hypergraph-based, and text-attributed hypergraph (TAHG)-based paradigms.
@@
Graph-based SSL. In this work, we adopt three widely used graph-based SSL frameworks: two contrastive learning methods, GraphCL You et al. (2020) and BGRL Thakoor et al. (2022), and one generative method, GraphMAE Hou et al. (2022). To adapt these methods to hypergraphs, we apply clique expansion following the protocol in Lee and Shin (2023). For textual feature extraction, we employ shallow encoders (SE) implemented via Skip-Gram Mikolov et al. (2013), as well as GIANT Chien et al. (2022a), a graph-agnostic text encoder designed for raw textual attributes.
@@
Hypergraph-based SSL. We consider four state-of-the-art self-supervised learning methods specifically designed for hypergraphs. These include two contrastive approaches, TriCL Lee and Shin (2023) and SE-HSSL Li et al. (2024); one generative framework, HypeBoy Kim et al. (2024a); and VilLain Lee et al. (2024), which is tailored for feature-absent hypergraph settings. For a fair comparison, all methods adopt the same shallow encoder (SE). In addition, we further evaluate these baselines by replacing the SE with two pre-trained language models, BERT-Base Devlin et al. (2019) and RoBERTa-Base Liu et al. (2019).
TAHG-based SSL. We further include HyperBERT Bazaga et al. (2024), to the best of our knowledge, the only existing SSL method specifically designed for text-attributed hypergraphs. HyperBERT jointly models textual semantics and hypergraph structure using a Transformer-based architecture, in which hypergraph-aware layers are introduced to integrate structural signals with text representations. However, this joint modeling paradigm incurs substantial computational and memory overhead, limiting its scalability to large hypergraphs, as validated in our experiments.
C.3 Evaluation Protocol
@@
We evaluate our method on two standard hypergraph learning tasks: node classification and hyperedge prediction. After pretraining with HiTeC, we adopt a linear evaluation protocol Lee and Shin (2023); Li et al. (2024); Kim et al. (2024a), where the pretrained node embeddings are frozen and used as input features for downstream classifiers.
@@
Node classification. Following Lee and Shin (2023); Li et al. (2024), we randomly split each dataset into 10%/10%/80% for training, validation, and testing. Then, a simple linear classifier is trained on top of the fixed node embeddings using $\ell_{2}$-regularized logistic regression. We report the average test accuracy and standard deviation over 20 random splits, each with 5 different initializations.
@@
Hyperedge prediction. Following Kim et al. (2024a), we need to generate negative hyperedge samples to evaluate a model on the hyperedge prediction task. In our experiments, we adopt Clique Negative Sampling (CNS) Yu et al. (2025), a widely used strategy for constructing negative hyperedge samples. Specifically, for a given hyperedge $e$, we replace a node $u\in e$ with a node $v\notin e$ such that $v$ is connected to all other nodes in $e$, i.e., $(e\setminus\{u\})\cup\{v\}$. Then, a two-layer MLP is trained on the fixed node embeddings using the 60%/20%/20% data split. Same as the node classification task, we report the average test accuracy and standard deviation over 20 random splits, each with 5 initializations.
C.4 Implementation Details
@@
All experiments are conducted on a single NVIDIA A800 PCIe GPU with 80GB of memory. For all baselines, we adopt their official implementations with default settings. For HiTeC, we adopt BERT-Base Devlin et al. (2019) as the default text encoder and apply LoRA Hu et al. (2022) to the query and value matrices of the self-attention layers for parameter-efficient fine-tuning, while keeping the PLM backbone frozen. Unless otherwise specified, the LoRA configuration is set to $r=16$, $\alpha=32$, and dropout =0.05=0.05. We additionally evaluate other text encoder backbones, including DistilBERT Sanh et al. (2019), RoBERTa-Base Liu et al. (2019), and DeBERTa-Base He et al. (2020). The text encoder is trained for 5 epochs on small datasets (Citeseer, Cora), 4 epochs on medium datasets (History, Photo), and 3 epochs on large datasets (Computers, Fitness), using a learning rate of 2e-5. The hypergraph encoder is implemented as a single-layer module. We tune the contrastive loss weights $\lambda_{e}$ and $\lambda_{s}$ over $\{1,2,3,4\}$. The overlap threshold $s$ and walk length $l$ are tuned from $\{1,2,3,4,5,10\}$, the subgraph sampling ratio $r$ in $\{10,20,30,40,50\}$. To ensure a fair comparison, we tune the learning rate from $\{0.01,0.001,1e\text{-}4,5e\text{-}4\}$ and weight decay from $\{0,1e\text{-}5\}$ for all methods.
Appendix D Additional Experiment Results